# Efficient high-resolution refinement in cryo-EM with stochastic gradient descent

### v2

@@ Abstract
Electron cryomicroscopy (cryo-EM) is an imaging technique widely used in structural biology to determine the three-dimensional structure of biological molecules from noisy two-dimensional projections with unknown orientations. As the typical pipeline involves processing large amounts of data, efficient algorithms are crucial for fast and reliable results. The stochastic gradient descent (SGD) algorithm has been used to improve the speed of ab initio reconstruction, which results in a first, low-resolution estimation of the volume representing the molecule of interest, but has yet to be applied successfully in the high-resolution regime, where expectation-maximization algorithms achieve state-of-the-art results, at a high computational cost. In this article, we investigate the conditioning of the optimization problem and show that the large condition number prevents the successful application of gradient descent-based methods at high resolution. Our results include a theoretical analysis of the condition number of the optimization problem in a simplified setting where the individual projection directions are known, an algorithm based on computing a diagonal preconditioner using Hutchinson’s diagonal estimator, and numerical experiments showing the improvement in the convergence speed when using the estimated preconditioner with SGD. The preconditioned SGD approach can potentially enable a simple and unified approach to ab initio reconstruction and high-resolution refinement with faster convergence speed and higher flexibility, and our results are a promising step in this direction.
@@ 1 Introduction
We consider the problem of gradient-based optimization for tomographic reconstruction with a particular focus on electron cryomicroscopy (cryo-EM). Stochastic gradient descent (SGD) based optimization methods have become a standard algorithmic framework in many areas due to their speed, robustness, flexibility and ease of implementation, particularly with the availability of fast and robust automatic differentiation libraries. However, in cryo-EM, SGD methods have only really found use in ab initio structure determination, where the goal is only a low-resolution structure. High-resolution structures are then determined by switching to a different form of optimization, typically some form of expectation-maximization (EM), despite the fact that these methods are often slow and require specific modeling assumptions. A natural question which we address here then is why stochastic gradient optimization techniques have not been able to solve the high-resolution optimization problem. Doing so could further speed up data processing for cryo-EM, simplify workflows and unify open research questions. However, perhaps more significantly, using SGD-based optimization methods would allow for more flexibility in modeling the reconstruction problem. Common modeling assumptions (e.g., Gaussian noise, Gaussian priors, rigid structures, discrete Fourier-based structure representations) which may not be optimal but are required by current refinement methods could be relaxed with SGD methods.
@@ 1.1 Background
Recently, a new class of methods for cryo-EM have emerged which aim to reconstruct not just static structures but also conformational heterogeneity, where the reconstructed volume is subject to different kinds of deformations [toader_methods_2023, lederman_hyper-molecules_2020, zhong_cryodrgn_2021, zhong_cryodrgn2_2021, levy_amortized_2022, chen_deep_2021, chen_integrating_2023, chen_improving_2023, punjani_3dva_2021, punjani_3dflex_2023, gupta_cryogan_2021, gupta_multi-cryogan_2020, herreros_zart_2023, gilles_bayesian_2023, li_cryostar_2023, schwab_dynamight_2023, esteve-yague_spectral_2023]. Such methods greatly expand the capabilities of cryo-EM, e.g., with time-resolved cyro-EM [torino_time-resolved_2023, bhattacharjee_mechanism_2023]. However, existing methods usually require a high resolution structure and known pose variables as input, limiting their applicability. Moreover, they cannot use the standard EM-based optimization approaches, often using neural networks trained using SGD instead, and generally fail to match existing rigid structure refinement approaches in resolution and requiring substantially more computation. While motivated by the promise of capturing conformational heterogeneity, we focus here on the static reconstruction case. Our results suggest that improving the performance of SGD-based methods may be the key to unlocking this new capability of cryo-EM.
1.2 Prior work
@@ 2.1 Forward model
Let $\phi_{i}\in SO(3)\times\operatorname{\mathbb{C}}^{2}$ denote the pose variable encapsulating the rotation matrix $R$ and the diagonal shift matrix $T$, and $\{r_{j}\}_{j=1}^{M_{x}}$ be the coordinates of the Fourier grid at $z=0$. We define the following projection operators $P_{\phi}\in\operatorname{\mathbb{C}}^{M_{x}\times M_{v}}$:
@@
    $(P^{nn}_{\phi}v)(r_{j}):=Tv_{k^{*}(j)},$ (1)
    where $k^{*}(j)=\operatorname*{argmin}_{k\in\{1,\ldots,M_{v}\}}\|r_{k}-R^{T}r_{j}\|_{2}$, for all $j=1,\ldots,M_{x}$. In this case, the matrix $P^{nn}_{\phi}$ has exactly one non-zero element in each row, and therefore ${(P^{nn}_{\phi})}^{*}P^{nn}_{\phi}\in\operatorname{\mathbb{R}}^{M_{v}\times M_{v}}$ is diagonal.
@@
    $(P^{tri}_{\phi}v)(r_{j}):=Tv_{j}^{*},$ (2)
    where $v_{j}^{*}$ is obtained by using trilinear interpolation of $v$ at $R^{T}r_{j}$, for all $j=1,\ldots,M_{x}$.
@@
An alternative approach to the interpolation given in Definition 2.1 is to sample the volume on the rotated grid using non-uniform FFT [dutt_fast_1993, greengard_accelerating_2004, barnett_parallel_2019] as done, for example, in [wang_fourier-based_2013]. In [anden_structural_2018], the structure of the matrix involving a projection and a backprojection is leveraged to obtain a fast preconditioner based on a circular convolution [tyrtyshnikov_optimal_1992].
@@ 2.3 Stochastic gradient descent for ab initio reconstruction
The SGD algorithm has two major advantages over EM. First, it has a lower computational cost per iteration, as it only uses a subset of the dataset, while EM requires a pass through the entire dataset. While computing the integrals in (5) for the images in the mini-batch cannot be avoided, the gridding and searching approaches that EM uses to efficiently sample the space of poses are also beneficial in the implementation of SGD. Second, it has been observed in practice that the noise in the sampled gradient enables SGD to explore the optimization landscape more efficiently before converging to a local minimizer of the objective function, preventing it from becoming stuck in unwanted local minima. Given these advantages, the SGD algorithm is particularly well-suited for ab initio reconstruction, as it has been shown in practice in cryoSPARC [punjani_cryosparc_2017]. More recently, the gradient-based algorithm VDAM has been introduced for ab initio reconstruction in the RELION software [kimanius_new_2021]. However, despite the clear benefits in terms of computational cost and convergence speed of stochastic gradient algorithms compared to EM for (low resolution) ab initio reconstruction, EM is still the state-of-the-art approach for high-resolution refinement.
@@
Note that we chose this particular result for its simplicity, despite the rather strong condition (9). More general results for the convergence of SGD where this condition is relaxed can be found, for example, in [moulines_non-asymptotic_2011, needell_stochastic_2016, gower_sgd_2019].
To improve the convergence of SGD, we will precondition it using an approximation of the diagonal of the Hessian matrix, which we will compute using Hutchinson’s diagonal estimator [bekas_estimator_2007]:
@@
where $z\in\mathbb{R}^{M_{v}}$ is a vector with elements drawn from a Rademacher(0.5) distribution (the entries of $z$ are $1$ or $-1$ with equal probability) and $\odot$ denotes element-wise multiplication.
The central claim of this article is that the condition number of the Hessian of the loss function in the cryo-EM reconstruction problem scales unfavorably with the target resolution of the reconstruction, and therefore it affects the convergence of SGD for high resolution refinement. In the next section, we argue that this is indeed the case and we propose a solution for overcoming this issue in a simplified setting.
@@ 3.1 Condition number
We will now analyze the structure and the condition number of the Hessian matrix $H$ when the projection matrices correspond to nearest-neighbor interpolation, namely $P_{i}=P_{\phi_{i}^{*}}^{nn}$, for $i=1,\ldots,N$. In order to do so, we introduce two necessary concepts. First, the projection assignment function of a particle image maps each element of the image to the element of the volume whose value is assigned to it by the projection operator.
@@
$v^{(k+1)}=v^{(k)}-\eta_{k}D^{-1}\nabla f(v^{(k)}).$ (22)
Remark 3.5.
For fixed $\sigma,\tau$, taking $\lambda=\sigma^{2}/\tau^{2}$ and known poses, the matrix whose inverse appears in the M step of the EM algorithm in equation (6b) is the full Hessian matrix (15) of the loss function. Therefore, EM implicitly solves the conditioning issue that is problematic for SGD, and in our preconditioning approach, we aim to approximate, using mini-batches, this gradient scaling that EM applies at every iteration.
@@ 3.2 Computing the preconditioner
The update of the preconditioner at the current iteration obtained using the current mini-batch is combined with the estimated preconditioner from the previous iteration using an exponential average as done, for example, in Adam [kingma_adam_2017], AdaHessian [yao_adahessian_2021] and OASIS [jahani_doubly_2022]. In addition, to take advantage of the fact that the Hessian $H$ in (15) of the objective function (12) is independent of the current iterate when the orientations are known, the exponential average is taken between the estimated preconditioner at the previous iteration and the estimated diagonal using all the samples of Rademacher vectors $z$ drawn up to the current iteration. However, in the more general reconstruction problem with unknown pose variables (4), only the current update would be used. Starting with the identity matrix as the initial estimate, $D^{(0)}=I_{M_{v}\times M_{v}}$, the update rule for the diagonal estimator $D^{(k)}$ at iteration $k$ is:
@@ 3.4 The algorithm
With the preconditioner estimation approach proposed in Section 3.2 and the thresholding described in Section 3.3, we present the full algorithm in Algorithm 1. No warm start is required for the approximation of the diagonal preconditioner, which is initialized with the identity matrix and estimated iteratively. The step size $\eta^{(k)}$ at each iteration is set using the stochastic Armijo line-search [vaswani_adaptive_2021], namely the largest step size $\eta$ is sought that satisfies the condition:
@@ 4 Numerical experiments
The dataset used in this section is derived from the Electron Microscopy Public Image Archive database entry EMPIAR-10076 of the bacterial ribosome, where $N=30000$ images of $192\times 192$ pixels are selected and whose pose variables have been computed using RELION333 The parameters were first obtained from the CryoDRGN tutorial [zhong_cryodrgn_2021], converted to RELION format using Pyem [asarnow_pyem_2019] and then further processed in RELION [scheres_relion_2012]. . The condition number of the reconstruction problem for this dataset with the computed pose variables is shown in Figure 3(a) for increasing values of the maximum Fourier radius $R$. Nearest-neighbor interpolation has been used for this figure, and therefore the condition number has been computed using (16). The condition number grows quickly with the resolution, and at the maximum radius $R$ corresponding to the grid side length $M=192$, it is of the order of $10^{2}$. The bound in Proposition (3.4) and equations (20)-(21), shown as the solid blue line in Figure 3(a), holds for the condition number when no CTF is included, shown as the red, dashed line. Because the CTF has magnitude less than one at the origin and the CTF envelope approaches zero at high spatial frequency, the numerator in (16) is smaller that in the case when no CTF is used, and therefore the condition number computed in this setting, shown in the green dashdot line, is not always larger than the theoretical lower bound. The effect of the CTF can be seen in the cross-section of the diagonal of the Hessian matrix $H$ in Figure 3(b).
@@
To verify the two claims on the dataset described above, we run three algorithms: SGD with no preconditioner, SGD with a preconditioner that has been precomputed in advance using (23), and Algorithm 1, namely SGD with a preconditioner estimated during the refinement process. All algorithms use the stochastic Armijo line-search (26) for the step size adaptation. While the theory in Section 3.1 applies to nearest-neighbor interpolation, where the matrices $P_{i}^{*}P_{i}$ are diagonal (and therefore the optimization problem can be solved easily by other methods), in the numerical experiments presented in this section, the projection operators are implemented using trilinear interpolation, which is often used in practice. In this case, the Hessian matrix $H$, whose diagonal we estimate and use as a preconditioner, is no longer diagonal.
@@
To evaluate and compare the performance of the three algorithms, we compute the value of the loss function in (12) after each epoch (a full pass through the dataset), as well as the Fourier Shell Correlation (FSC) of each reconstruction with a ground-truth solution, obtained using L-BFGS with 1000 iterations. The FSC, a standard error measure in the cryo-EM literature, is the cross-correlation coefficient between two volumes across three-dimensional shells in the Fourier domain [harauz_exact_1986]. Specifically, given two volumes $u$ and $v$ in the Fourier space, the FSC at radius $r$ from the origin is defined as
@@
where $S_{r}$ is the set of Fourier voxels in the spherical shell at radius $r$.
Figure 4 shows the results obtained using the three algorithms. In panel (a), we see the convergence of the loss function significantly improved when using the precomputed preconditioner (red) over not using a preconditioner (blue), verifying Claim 1 above. When estimating the preconditioner during refinement using Algorithm 1, the value of the loss (green, solid) is between those of the previous two algorithms and approaches the fully preconditioned SGD as the preconditioner progressively becomes more accurate, verifying Claim 2. In panel (b), we show the FSC between the final reconstructions of each of the three algorithms and the ground truth reconstruction. While at low resolution, the FSC value is high for all algorithms, showing good convergence, at high resolution the FSC degrades for SGD with no preconditioning (blue), while the two preconditioned algorithms have a higher value, closer to one.

### v3

@@ Abstract
Electron cryo-microscopy (cryo-EM) is an imaging technique widely used in structural biology to determine the three-dimensional structure of biological molecules from noisy two-dimensional projections with unknown orientations. As the typical pipeline involves processing large amounts of data, efficient algorithms are crucial for fast and reliable results. The stochastic gradient descent (SGD) algorithm has been used to improve the speed of ab initio reconstruction, which results in a first, low-resolution estimation of the volume representing the molecule of interest, but has yet to be applied successfully in the high-resolution regime, where expectation-maximization algorithms achieve state-of-the-art results, at a high computational cost. In this article, we investigate the conditioning of the optimization problem and show that the large condition number prevents the successful application of gradient descent-based methods at high resolution. Our results include a theoretical analysis of the condition number of the optimization problem in a simplified setting where the individual projection directions are known, an algorithm based on computing a diagonal preconditioner using Hutchinson’s diagonal estimator, and numerical experiments showing the improvement in the convergence speed when using the estimated preconditioner with SGD. The preconditioned SGD approach can potentially enable a simple and unified approach to ab initio reconstruction and high-resolution refinement with faster convergence speed and higher flexibility, and our results are a promising step in this direction.
@@ 1 Introduction
We consider the problem of gradient-based optimization for tomographic reconstruction with a particular focus on electron cryo-microscopy (cryo-EM). Stochastic gradient descent (SGD) based optimization methods have become a standard algorithmic framework in many areas due to their speed, robustness, flexibility and ease of implementation, particularly with the availability of fast and robust automatic differentiation libraries. However, in cryo-EM, SGD methods have only really found use in ab initio structure determination, where the goal is only a low-resolution structure. High-resolution structures are then determined by switching to a different form of optimization, typically some form of expectation-maximization (EM), despite the fact that these methods are often slow and require specific modeling assumptions. A natural question which we address here then is why stochastic gradient optimization techniques have not been able to solve the high-resolution optimization problem. Doing so could further speed up data processing for cryo-EM, simplify workflows and unify open research questions. However, perhaps more significantly, using SGD-based optimization methods would allow for more flexibility in modeling the reconstruction problem. Common modeling assumptions (e.g., Gaussian noise, Gaussian priors, rigid structures, discrete Fourier-based structure representations) which may not be optimal but are required by current refinement methods could be relaxed with SGD methods.
@@ 1.1 Background
Recently, a new class of methods for cryo-EM have emerged which aim to reconstruct not just static structures but also conformational heterogeneity, where the reconstructed volume is subject to different kinds of deformations [toader_methods_2023, lederman_hyper-molecules_2020, zhong_cryodrgn_2021, zhong_cryodrgn2_2021, levy_amortized_2022, chen_deep_2021, chen_integrating_2023, chen_improving_2023, punjani_3dva_2021, punjani_3dflex_2023, gupta_cryogan_2021, gupta_multi-cryogan_2020, herreros_zart_2023, esteve-yague_spectral_2023, schwab_dynamight_2024, li_cryostar_2024, gilles_cryo-em_2025, woollard_instamap_2025]. Such methods greatly expand the capabilities of cryo-EM, e.g., with time-resolved cyro-EM [torino_time-resolved_2023, bhattacharjee_time_2024]. However, existing methods usually require a high resolution structure and known pose variables as input, limiting their applicability. Moreover, they cannot use the standard EM-based optimization approaches, often using neural networks trained using SGD instead, and generally fail to match existing rigid structure refinement approaches in resolution and requiring substantially more computation. While motivated by the promise of capturing conformational heterogeneity, we focus here on the static reconstruction case. Our results suggest that improving the performance of SGD-based methods may be the key to unlocking this new capability of cryo-EM.
1.2 Comparison to prior work
@@ 2.1 Forward model
Let $\phi_{i}\in SO(3)\times\operatorname{\mathbb{C}}^{2}$ denote the pose variable encapsulating the rotation matrix $R\in\operatorname{\mathbb{R}}^{3\times 3}$ and the diagonal shift matrix $T\in\operatorname{\mathbb{C}}^{M_{x}\times M_{x}}$, $\{r_{j}^{2D}\}_{j=1}^{M_{x}}$ the coordinates of the Fourier grid at the $z=0$ plane, and $\{r_{k}^{3D}\}_{k=1}^{M_{v}}$ the coordinates of the 3D Fourier grid that the volume $v$ is defined on. We define the following projection operators $P_{\phi}\in\operatorname{\mathbb{C}}^{M_{x}\times M_{v}}$:
@@
    $(P^{nn}_{\phi}v)_{j}:=T_{jj}v_{k^{*}(j)},\qquad\text{for all}\quad j=1,\ldots,M_{x},$ (1)
    where $k^{*}(j)=\operatorname*{argmin}_{k\in\{1,\ldots,M_{v}\}}\|r_{k}^{3D}-R^{T}r_{j}^{2D}\|_{2}$, In this case, the matrix $P^{nn}_{\phi}$ has exactly one non-zero element in each row, and therefore ${(P^{nn}_{\phi})}^{*}P^{nn}_{\phi}\in\operatorname{\mathbb{R}}^{M_{v}\times M_{v}}$ is diagonal.
@@
    $(P^{tri}_{\phi}v)_{j}:=T_{jj}\sum_{i=1}^{8}c_{k_{i}(j)}v_{k_{i}(j)},\qquad\text{for all}\quad j=1,\ldots,M_{x}.$ (2)
    where $k_{1}(j),\ldots,k_{8}(j)$ are the indices of the eight closest 3D grid points to the rotated 2D grid point $R^{T}r_{j}^{2D}$ and $c_{k_{1}(j)},\ldots,c_{k_{8}(j)}$ are the weights of the trilinear interpolation of the volume $v$ at $R^{T}r_{j}^{2D}$ on these eight grid points.
@@
An alternative approach to the interpolation given in Definition 2.1 is to sample the volume on the rotated grid using non-uniform FFT [dutt_fast_1993, greengard_accelerating_2004, barnett_parallel_2019] as done, for example, in [wang_fourier-based_2013]. In [anden_structural_2018], the structure of the matrix involving a projection and a backprojection is leveraged to obtain a fast preconditioner based on a circular convolution [tyrtyshnikov_optimal_1992]. However, this setup is less suitable for our problem where the goal is to use stochastic gradient descent to solve the general volume reconstruction problem with unknown poses.
@@ 2.3 Stochastic gradient descent for ab initio reconstruction
The SGD algorithm has two major advantages over EM. First, it has a lower computational cost per iteration, as it only uses a subset of the dataset, while EM requires a pass through the entire dataset. While computing the integrals in (5) for the images in the mini-batch cannot be avoided, the gridding and searching approaches that EM uses to efficiently sample the space of poses are also beneficial in the implementation of SGD. Second, it has been observed that the noise in the sampled gradient enables SGD to explore the optimization landscape more efficiently, preventing it from becoming stuck in unwanted local minima [keskar_large-batch_2017, smith_bayesian_2018]. Given these advantages, the SGD algorithm is particularly well-suited for ab initio reconstruction, as it has been shown in practice in cryoSPARC [punjani_cryosparc_2017]. More recently, the gradient-based algorithm VDAM has been introduced for ab initio reconstruction in the RELION software [kimanius_new_2021]. However, despite the clear benefits in terms of computational cost and convergence speed of stochastic gradient algorithms compared to EM for (low resolution) ab initio reconstruction, EM is still the state-of-the-art approach for high-resolution refinement.
@@
Put informally, this means that the number of iterations required to reach a certain accuracy $\epsilon$ scales with the condition number of the Hessian or, in other words, the impact of a condition number equal to $C>1$ is that $C$ times more iterations are required to reach the same accuracy $\epsilon$ compared to the case where the condition number is 1. Note that we chose this particular result for its simplicity, despite the rather strong condition (9). More general results for the convergence of SGD where this condition is relaxed can be found, for example, in [moulines_non-asymptotic_2011, needell_stochastic_2016, gower_sgd_2019].
In general, one can apply a preconditioner to improve the convergence of SGD. Preconditioning linearly reparameterizes the optimization problem to change its condition number without changing its solution in order to improve numerical behaviour and convergence dynamics. In particular, we will precondition the optimization problem using an approximation of the diagonal of the Hessian matrix of the loss function, which we will define precisely in the next section. Such an approximation of the Hessian diagonal will be obtained using Hutchinson’s diagonal estimator [bekas_estimator_2007], which allows one to compute (an approximation of) the diagonal of a matrix $H$ only using matrix-vector products, without forming the full matrix $H$. This is achieved by computing the expectation:
@@
where $z\in\mathbb{R}^{M_{v}}$ is a vector with elements drawn from a Rademacher(0.5) distribution (the elements of $z$ are $1$ or $-1$ with equal probability) and $\odot$ denotes element-wise multiplication.
The central claim of this article is that the condition number of the Hessian of the loss function in the cryo-EM reconstruction problem scales with the target resolution of the reconstruction, and slows the convergence of SGD for high resolution refinement. In the next section, we argue that this is indeed the case and we propose a solution for overcoming this issue in a simplified setting.
@@ 3.1 Condition number
We will now analyze the structure and the condition number of the Hessian matrix $H$ when the projection matrices correspond to nearest-neighbor interpolation, namely $P_{i}=P_{\phi_{i}^{*}}^{nn}$, for $i=1,\ldots,N$, as given in Definition 2.1. In order to do so, we introduce two necessary concepts. First, the projection assignment function of a particle image maps each element of the image to the element of the volume whose value is assigned to it by the projection operator.
@@
$v^{(k+1)}=v^{(k)}-\eta_{k}D^{-1}\nabla f(v^{(k)}).$ (22)
Remark 3.6.
For fixed $\sigma,\tau$, taking $\lambda=\sigma^{2}/\tau^{2}$ and known poses, the matrix whose inverse appears in the M step of the EM algorithm in equation (6b) is the full Hessian matrix $H$ in (15) of the loss function, and so the M step becomes:
$v^{(k+1)}=\left(\sum_{i=1}^{N}P_{i}^{*}P_{i}+\lambda I_{M_{v}}\right)^{-1}\left(\sum_{i=1}^{N}P_{i}^{*}x_{i}\right)=H^{-1}\left(\sum_{i=1}^{N}P_{i}^{*}x_{i},\right)$
Therefore, EM implicitly solves the conditioning issue that is problematic for SGD, and in our preconditioning approach, we aim to approximate, using mini-batches, the diagonal part of this gradient scaling that EM applies at every iteration.
@@ 3.2 Computing the preconditioner
The update of the preconditioner at the current iteration obtained using the current mini-batch is combined with the estimated preconditioner from the previous iteration using an exponential average as done, for example, in Adam [kingma_adam_2015], AdaHessian [yao_adahessian_2021] and OASIS [jahani_doubly_2022]. In addition, to take advantage of the fact that the Hessian $H$ in (15) of the objective function (12) is independent of the current iterate when the orientations are known, the exponential average is taken between the estimated preconditioner at the previous iteration and the estimated diagonal using all the samples of Rademacher vectors $z$ drawn up to the current iteration. However, in the more general reconstruction problem with unknown pose variables (4), only the current update would be used. Starting with the identity matrix as the initial estimate, $D^{(0)}=I_{M_{v}\times M_{v}}$, the update rule for the diagonal estimator $D^{(k)}$ at iteration $k$ is:
@@ 3.4 The algorithm
With the preconditioner estimation approach proposed in Section 3.2 and the thresholding described in Section 3.3, we present the full algorithm in Algorithm 1. No warm start is required for the approximation of the diagonal preconditioner, which is initialized with the identity matrix $I_{n}$ and estimated iteratively; the running average $D_{avg}$ computed as part of the diagonal approximation (23) is initialized as the zero $n\times n$ matrix $O_{n}$. The step size $\eta^{(k)}$ at each iteration is set using the stochastic Armijo line-search [vaswani_adaptive_2021], namely the largest step size $\eta$ is sought that satisfies the condition:
@@ 4.1 Setup
The dataset used in this section is derived from the Electron Microscopy Public Image Archive database entry EMPIAR-10076 of the bacterial ribosome, where $N=30000$ images of $192\times 192$ pixels are selected and whose pose variables have been computed using RELION333 The parameters were first obtained from the CryoDRGN tutorial [zhong_cryodrgn_2021], converted to RELION format using Pyem [asarnow_pyem_2019] and then further processed in RELION [scheres_relion_2012]. Specifically, one round of 3D refinement and one round of 3D classification with 8 classes were performed, followed by one final 3D refinement of one of the resulting homogeneous classes, consisting of $N=33026$ particles, to a resolution of $4.2$. $N=30000$ particles were selected at random from the resulting homogeneous class. While the complete removal of all continuous heterogeneity is not relevant for this paper and therefore has not been attempted, the resulting dataset is still a good approximation of the EMD-8455 and EMD-8456 discrete classes associated with the EMPIAR-10076 entry. . The condition number of the reconstruction problem for this dataset with the computed pose variables is shown in Figure 3(a) for increasing values of the maximum Fourier radius $R$. Nearest-neighbor interpolation has been used for this figure, and therefore the condition number has been computed using (16). The condition number grows quickly with the resolution, and at the maximum radius $R$ corresponding to the grid side length $M=192$, it is of the order of $10^{2}$. The bound in Proposition (3.4) and equations (20)-(21), shown as the solid blue line in Figure 3(a), holds for the condition number when no CTF is included, shown as the red, dashed line. While the theory in Section 3 does not handle the case when CTF is used, we will show empirically in Section 4.2 that the same trend of the ill-conditionning increasing with the resolution holds when the particle images, extracted from true micrographs, are corrupted by the CTF. To see why the theory does not apply, note that because the CTF has magnitude less than one at the origin and it oscillates around zero, the numerator in (16) is smaller than in the case when no CTF is used, and therefore the condition number computed in this setting, shown in the green dashdot line, is not always larger than the theoretical lower bound. Figure 3(b), showing a one-dimensional cross-section through the diagonal of the Hessian both with and without CTF in the same setting as in Figure 3(a), explains the shape of the green curve in Figure 3(a): both the minimum and the maximum values of $\text{diag}(H)$ occur in the interval $[0,20]$ (in voxels). The effect of the CTF is further illustrated in Figure 3(c), in the two-dimensional cross-section of the diagonal of the Hessian matrix $H$ in the case when trilinear interpolation is used. While the CTF leads to a Hessian condition number that is lower than the one without CTF, the difficulty of the underlying inverse problem is not decreased: as the CTF removes information from the images, the difficulty of the reconstruction problem is actually increased.
@@
To verify the two claims on the dataset described above, we run three algorithms: SGD with no preconditioner; SGD with a preconditioner that has been precomputed in advance using (23) over 1000 epochs with batch size 1000; and Algorithm 1, namely SGD with a preconditioner estimated during the refinement process over 10 epochs. All algorithms are initialized with the same random volume obtained by sampling each voxel value from a complex normal distribution, and use the stochastic Armijo line-search (26) for the step size adaptation. While the theory in Section 3.1 applies to nearest-neighbor interpolation, where the matrices $P_{i}^{*}P_{i}$ are diagonal (and therefore the optimization problem can be solved easily by other methods), in the numerical experiments presented in this section, the projection operators are implemented using trilinear interpolation, which is often used in practice. In this case, the Hessian matrix $H$, whose diagonal we estimate and use as a preconditioner, is no longer diagonal.
Remark 4.1.
The convergence of Algorithm 1 is expected to be at most as good as the convergence of SGD with the precomputed preconditioner. This is due to the fact that the preconditioner has been computed using 1000 epochs of (23) and this preconditioner is then used from the start of the SGD algorithm, while Algorithm 1 starts with the identity matrix as the preconditioner and estimates it during the 10 epochs (using (23) as well). If the particle poses are known and fixed, one can simply precompute the preconditioner corresponding to the forward operator with the given poses in advance. However, the aim of the proposed method for estimating the preconditioner is to be incorporated into a reconstruction algorithm that estimates the poses in addition to the volume. By doing so, the optimal preconditioner changes when pose variables are updated, making it impossible to precompute a good preconditioner in advance. Our aim is to show that over only a small number of epochs (10 in our numerical experiments), the proposed method computes a good enough preconditioner that allows obtaining a solution to the inverse problem at high-frequencies.
To evaluate and compare the performance of the three algorithms, we first compute a numerical ground truth solution to the optimization problem (12) using the L-BFGS algorithm [liu_limited_1989], which we run for 1000 iterations. This approach to obtaining a ground truth benefits from the convexity of problem (12), as well as the good convergence properties of the L-BFGS algorithm due to its estimation of the Hessian of the loss function. By running it for a number of iterations much larger than the number of epochs for which we would run SGD in a practical setting in cryo-EM reconstruction, we ensure that the L-BFGS solution is a good baseline against which to evaluate the SGD solutions. We show the convergence of L-BFGS on the given dataset and problem (12) in Figure 3(d).
@@
In the runs of the three SGD algorithms, we compute the value of the loss function in (12) after each epoch (a full pass through the dataset), as well as the Fourier Shell Correlation (FSC) of each reconstruction with the L-BFGS ground truth solution.
The FSC, a standard error measure in the cryo-EM literature, is the cross-correlation coefficient between two volumes across three-dimensional shells in the Fourier domain [harauz_exact_1986]. Specifically, given two volumes $u$ and $v$ in the Fourier space, the FSC at radius $r$ from the origin is defined as
@@
where $S_{r}$ is the set of Fourier voxels in the spherical shell at radius $r$. While the FSC is used in the cryo-EM literature as a measure of the resolution or consistency of a reconstruction, here it will be used differently. Instead, we will compute the FSC between the iterates of each SGD algorithm and the L-BFGS ground truth volume in order to understand how different frequency bands of the structures are converging at different rates.
4.2 Results
Figure 4 shows the results obtained using the three algorithms. In panel (a), we see the convergence of the loss function significantly improved when using the precomputed preconditioner (red) over not using a preconditioner (blue), verifying Claim 1 above. When estimating the preconditioner during refinement using Algorithm 1, the value of the loss (green, solid) is between those of the previous two algorithms and approaches the fully preconditioned SGD as the preconditioner progressively becomes more accurate, verifying Claim 2. For reference, the black dashed line shows the loss function value at the final volume obtained using L-BFGS and used as a ground truth reconstruction. As expected, this is lower than all of the three SGD algorithms that we evaluate in this section. In panel (b), we show the FSC between the final reconstructions of each of the three algorithms and the ground truth reconstruction. While at low resolution, the FSC value is high for all algorithms, showing good convergence, at high resolution the FSC degrades for SGD with no preconditioning (blue), while the two preconditioned algorithms have a higher value, closer to one.